A gumball machine contains 300 grape flavored balls, 400 cherry flavored balls, and 500 lemon flavored balls. What is the probability of getting 1 grape ball, 1 cherry ball, and 1 lemon ball if each ball was removed and then replaced before choosing the next from the machine?



1
Expert's answer
2021-12-02T18:38:49-0500

 Probability of getting one grape, 1 cherry and 1 lemon ball = 0.03472

Numbers of grapes flavored balls = 300

Numbers of cheer flavored balls = 400

Numbers of lemon flavored balls = 500

Total number of balls = 300 + 400 + 500 = 1200

Probability of getting 1 grape ball = 300/1200 = 0.25

Probability of getting 1 cherry ball = 400/1200 = 0.3334

Probability of getting 1 lemon ball = 500/1200 = 0.0.41667

Probability of getting one grape, 1 cherry and 1 lemon ball = "\frac{300}{1200}\times\frac{400}{1200}\times\frac{500}{1200}=0.03472"